OpenStudy (cutiecomittee123):

solve the equation for x 4000-1000e^2x

OpenStudy (anonymous):

want some help

OpenStudy (cutiecomittee123):

of course

OpenStudy (anonymous):

lol ok....so one trick to make this equation easier is to take all them zeros lol

OpenStudy (cutiecomittee123):

okay so make it 4=1e^2x

OpenStudy (anonymous):

ok then do the distributive property

OpenStudy (cutiecomittee123):

how so ?

OpenStudy (agl202):

4000=1000e^2x e^2x=4 take log both sides log(e^2x)=log4 2x*loge=log4 2x=log4 {bcoz loge=1} 2x=.6020 x=.6020/2=.3010 x=log4/2 is the correct one...

OpenStudy (anonymous):

2x*loge=log4 2x=log4 {bcoz loge=1} 2x=.6020 x=.6020/2=.3010

OpenStudy (cutiecomittee123):

thanks girls!

OpenStudy (anonymous):

np

OpenStudy (cutiecomittee123):

wait

OpenStudy (anonymous):

......

OpenStudy (cutiecomittee123):

the options for the solution are x=1n4-2 x=4/e^2 x=1n4/2 x=ln40000/2Xln10000

OpenStudy (cutiecomittee123):

so the answer id x=ln4/2?????

OpenStudy (cutiecomittee123):

is*